OpenStudy (oliviagggg):

What is the common ratio of the sequence 256,128,64,32,...

OpenStudy (freckles):

are you having trouble diving some term by it's previous term?

OpenStudy (oliviagggg):

I can't remember the pattern for the division

OpenStudy (freckles):

dividing*

OpenStudy (f_jayyy):

look at each number closely...they are all decreasing....by how much??? 256/2 = 128....128/2=64...so the common sequence is decreasing by half

OpenStudy (freckles):

\[\frac{\text{ choose a term } } { \text{ previous term goes here }} \]

OpenStudy (oliviagggg):

OK I get it now thank you the rest of my questions are like this so this is all I need

OpenStudy (freckles):

np